### v1

@@ 3.1.7 Constraining motif: the $Q$$X$ module
    The conditional division time distribution, conditioned on the initial size, and also the steady state division time distribution predicted by this model exactly match the corresponding experimental distribution, solving an outstanding puzzle that arose from mismatch between data and a size-thresholded division model of stochastic exponential growth [52, 37]!
@@ 4 CONCLUSION AND FUTURE PERSPECTIVE
In this review, we have illustrated how an integrated approach uniting high-precision measurements, data-informed insights, and physics theory can successfully address outstanding questions of fundamental biology. Focusing on the phenomenon of stochastic growth and division processes in the simplest living organism (the bacterial cell), we have walked through a procedure for analyzing high-throughput, high-precision dynamic datasets to identify emergent simplicities, in particular various scaling laws, that provide new insights into a long-standing problem (that of cell size homeostasis). Recasting the question from a stochastic, intergenerational viewpoint (i.e., one that considers the entire life histories of individual cells without recourse to a priori mechanistic assumptions), and taking advantage of identified emergent simplicities to achieve dimensional reduction of the problem, permits a reformulation that captures the inherent stochasticity of individual cells.
Identification of discrete modes by which homeostasis is maintained—in particular, via reflexive (elastic) adaptation of cell size and reflective (plastic) adaptation of growth rate—provides important insights into key system constraints that govern living bacterial cells, with additional implications for the design of functional adaptive synthetic homeostats. The observation of non-Markovian dynamics in single-cell growth rates implies the existence of intergenerational memory and plastic adaptation in these simple organisms, presenting an exciting opportunity to uncover the mechanistic basis of learned behavior in a single-celled life form without neurons [90, 91].

### v2

@@ 3.1.7 Constraining motif: the $Q$$X$ module
    The conditional division time distribution, conditioned on the initial size, and also the steady-state division time distribution predicted by this model exactly match the corresponding experimental distribution, solving an outstanding puzzle that arose from mismatch between data and a size-thresholded division model of stochastic exponential growth [52, 37]!
@@ 4 CONCLUSION AND FUTURE PERSPECTIVE
General principles of biology tend to be rare, owing to the idiosyncratic life histories of different species. An obvious exception is evolution by drift and selection, which provides a powerful framework encompassing the full diversity of life. Allometry describes systematic and regular scaling of organismal form and function corresponding to variations in body size, which arises from the connection between strong physical constraints and optimal physiology [90]. These interspecies “scaling” relationships represent a certain type of universality, based on shared dominant constraints. However, one of the challenges to a general theory of living systems has been that the sources of cross-species regularity tend to be statistical phenomenologies, rather than being derived from first principles. Bridging constitutive dynamics with macroscale regularity requires an approach that is both sensitive to genetic circuitry and also mappable onto low-dimensional, organism-level “state variables”, which provide the most efficient, and perhaps most predictive, descriptions of biology.
In this review, we have illustrated how an integrated approach uniting high-precision measurements, data-informed insights, and physics theory can successfully address outstanding questions of fundamental biology. Focusing on the phenomenon of stochastic growth and division processes in the simplest living organism (the bacterial cell), we have walked through a procedure for analyzing high-throughput, high-precision dynamic datasets to identify emergent simplicities, in particular various scaling laws, that provide new insights into a long-standing problem (that of cell size homeostasis). These scaling laws include the mean-rescaled cell size distributions across cell ages, the mean-rescaled distributions of division time and cell age across growth conditions, and the mean-rescaled distributions linking division ratio to both growth rate and division time ($re^{k\tau}$) across experimental modalities. Recasting the question from a stochastic, intergenerational viewpoint (i.e., one that considers the entire life histories of individual cells without recourse to a priori mechanistic assumptions), and taking advantage of identified emergent simplicities to achieve dimensional reduction of the problem, permits a reformulation that captures the inherent stochasticity of individual cells.
Identification of discrete modes by which homeostasis is maintained—in particular, via reflexive (elastic) adaptation of cell size and reflective (plastic) adaptation of growth rate—provides important insights into key system constraints that govern living bacterial cells, with additional implications for the design of functional adaptive synthetic homeostats. The observation of non-Markovian dynamics in single-cell growth rates implies the existence of intergenerational memory and plastic adaptation in these simple organisms, presenting an exciting opportunity to uncover the mechanistic basis of learned behavior in a single-celled life form without neurons [91, 92]. In particular, the generalization from steady state to complex, time-varying growth conditions presents many promising avenues for future research. This includes transitions between different growth regimes, growth under stressful conditions that lead to cell death and aging (such as starvation or exposure to antibiotics), and growth in environments that permit non-trivial collective behaviors emerging from interactions between cells.